OpenStudy (anonymous):

What is the volume of a square pyramid with base edges of 90 yards and a height of 20 yards? 36,000 yd3 162,000 yd3 54,000 yd3 1,800 yd3

Parth (parthkohli):

\[V = \dfrac{b \times b \times h }{3}\]

OpenStudy (nurali):

volume =1/3 x base area x height =1/3 x r^2 x h = 1/3 x 90^2 x 20 = 54000